OpenStudy (anonymous):

pls help will give medal A cylinder and a rectangular prism have the same volume and same height. The base of the prism is a square with side length 7 cm. What is the approximate radius of the cylinder? A. 4.0 cm B. 3.5 cm C. 2.2 cm D. 1.5 cm

OpenStudy (anonymous):

volume of...cylinder = pi r^2 h where h = height of the cylinder r = radius of the cylinder and volume of rectangular prism = l.b.h. where l = length b = breadth and h = height and since as it has been given that its a square with side length 7cm therefore l *b= 7*7

OpenStudy (anonymous):

7 x 7 is 49

OpenStudy (anonymous):

so the answer is a

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

wait someone is going to check it

OpenStudy (anonymous):

and np

OpenStudy (anonymous):

@davidmack

OpenStudy (anonymous):

V=pi*r^2*h for a cylinder V=l*w*h for a prism the volumes are equal pi*r^2*h=l*w*h the heights are equal, so divide both sides by h pi*r^2=l*w l and w are both 7 cm (base is a square with side lengths 7 cm) pi*r^2=7*7 pi*r^2=49 r^2=49/pi r^2=49/3.14159 r^2=15.597197 take the square root of both sides r=3.94932875 cm 3.94932875≈4, the answer is A) 4.0 cm

OpenStudy (anonymous):

thanks